Please help!
A collection of 40 dimes and nickels is worth $2.90. How many nickels are there? (hint: In your equation, use 290 cents instead of $2.90.)

Answer by checkley77(12844)

d+n=40
d=40-n
10d+5n=290
10(40-n)+5n=290
400-10n+5n=290
-5n=290-400
-5n=-110
n=-110/-5
n=22 nickles.
40-22=18 dimes.
Proof:
10*18+5*22=290
180+110=290
290=290